OpenStudy (anonymous):

Input the equation of the given line in standard form. The line with m =-1/2 and the point (-2, -3).

OpenStudy (anonymous):

Would really appreciate help

OpenStudy (xmoses1):

Use this Equation: \[y-y ^{1}=m(x-x ^{1})\]

OpenStudy (xmoses1):

I can write it out for you and then you can do the rest. \[y=3=-\frac{ 1 }{ 2 }x+1\] Now you work it out from there :)

OpenStudy (anonymous):

Thank you very much! (:

OpenStudy (xmoses1):

Your welcome! I'm glad I could help you out :)

OpenStudy (xmoses1):

And btw i meant to write y-3 not y=3 lol sorry